OpenStudy (anonymous):

solve the following equation -(3y+4)-6y=-3 (3y+2)

OpenStudy (anonymous):

-3y-4-6y=-9y-6 -9y=-9y-2 0=-2 no solution

OpenStudy (anonymous):

im pretty it's no solution!!!